OpenStudy (anonymous):

two angles are complements of each other the first angle has a measure of 17x+14 and other angle has a measure of 11x-8 solve for x

OpenStudy (anonymous):

Complementary angles means the sum of the two angles = 90 degrees SO: (First Angle) + (Second Angle) = 90 First Angle = 17x + 14 Second Angle = 11x - 8 Plug and chug you get: 17x + 14 + 11x - 8 = 90 degrees Solve for X

OpenStudy (anonymous):

so its 17x+11x=90-14+8

OpenStudy (anonymous):

Yup... take it a bit further and you get 28x = 84 then solving for x you get x = 3